# The infinitesimal structure of manifolds with non-continuous Riemannian metrics

### v1

@@ 3 The distance on manifolds with non-continuous Riemannian metrics
We may assume to be working in $\mathbb{R}^{d}$. Now there exist $\lambda,\Lambda>0$ such that $\lambda Id\leq g\leq\Lambda Id$ on $K$. Denote $v:=\dot{\gamma}^{1}_{t}$ and $w:=\dot{\gamma}^{2}_{t}$. By translation we may assume that $t=0$ and denote $p:=\gamma^{1}_{0}=\gamma^{2}_{0}$. By the previous proposition, we may assume that $\gamma^{1}_{t}=p+tv$ and $\gamma^{2}_{t}=p+tw$. But then
@@
Given a manifold $M$, a distance $\mathsf{d}$ on $M$, and a measurable positive definite inner product $g$ on $TM$, we call $(M,\mathsf{d},g)$ a quasi-Riemannian manifold if there exists an atlas $\mathcal{A}=\{(W,\psi)\}$ such that the following conditions hold:
@@
    In each $(W,\psi)$, there exist constants $\lambda,\Lambda>0$ such that for any $x,y\in\psi(W)$, it holds
    $\lambda\leq\sum_{i,j}(\psi_{*}g)^{ij}\partial_{x^{i}}\mathsf{d}(\psi^{-1}(x),\psi^{-1}(y))\partial_{x^{j}}\mathsf{d}(\psi^{-1}(x),\psi^{-1}(y))\leq\Lambda.$
@@
    $\limsup_{y\to x}\frac{|f(x)-f(y)|}{\mathsf{d}(x,y)}=|\nabla_{g}f|_{g}.$
@@ 4 Construction of examples
In this section, we will construct Riemannian manifolds with non-continuous metrics that together with the distance $\mathsf{d}_{g}$ as defined in the previous section, do not give rise to quasi-Riemannian manifolds or infinitesimally Hilbertian metric measure spaces.
4.1 A first example of a non-quasi Riemannian manifold
We will now construct a metric $g$ on $\mathbb{R}^{d}$ such that $g,g^{-1}\in L^{\infty}$ and such that there exists a smooth function $f:\mathbb{R}^{d}\to\mathbb{R}$ such that $|Df|_{w},|Df|>|\nabla_{g}f|_{g}$ on a set of positive measure.
@@ 4.1 A first example of a non-quasi Riemannian manifold
$g(x_{1},\ldots,x_{d}):=\theta(x_{1})Id\in L^{\infty}(\mathbb{R}^{d},\mathbb{R}^{d\times d}).$
Then $(\mathbb{R}^{d},g)$ satisfies (ii) from Assumption 3.1 and $(\mathbb{R}^{d},\mathsf{d}_{g},\mathrm{vol}_{g})$ satisfies condition (2.1). We observe that
@@
Proof of the claim. Fix $t,h$ as above and fix an $\varepsilon>0$. Note that from $g\geq Id$, we get that $\mathsf{d}_{g}\geq\mathsf{d}_{euc}$, hence $\mathsf{d}_{g}(\gamma^{y}_{t},\gamma^{y}_{t+h})\geq h$. Thus, it remains to prove that $\mathsf{d}_{g}(\gamma^{y}_{t},\gamma^{y}_{t+h})\leq h$. There exists a $q\in\mathbb{Q}\cap(0,1)$ such that $|x_{1}(y)-q|\leq\varepsilon$. Let $i\in\mathbb{N}$ be such that $q_{i}=q$ in the chosen counting. Then define the curves
@@
Denote by $\lambda$ the concatenation of those three curves and note that it yields a piecewise smooth curve that connects $\gamma^{y}_{t}$ and $\gamma^{y}_{t+h}$. Now, for all $\varsigma>0$, it holds $g_{\varsigma}=g*\rho_{\varsigma}\leq 2Id$, so we get that $L_{g_{\varsigma}}(\lambda_{1})+L_{g_{\varsigma}}(\lambda_{3})\leq 4\varepsilon$. For $\varsigma\leq\frac{\kappa}{2^{i+2}}$ that $g_{\varsigma}|_{\{x_{1}=q\}}\equiv Id$, hence $L_{g_{\varsigma}}(\lambda_{2})=h$. Thus,
@@
Note that this is indeed a feasible test plan, as ${\rm d}\mathcal{L}^{d}\leq{\rm d}\mathrm{vol}_{g}\leq 2^{\frac{d}{2}}{\rm d}\mathcal{L}^{d}$. But now for all $y\in E$, almost all $t\in[0,1]$, we get that $g(\gamma^{y}_{t})\geq\frac{3}{2}Id$ and $\dot{\gamma}^{y}_{t}=e_{2}$ and $|\dot{\gamma}^{y}_{t}|_{g}\geq\sqrt{\frac{3}{2}}$. But by (4.3), we get that the metric speed $|\dot{\gamma}^{y}_{t}|=1$ for all $y,t$, hence we found a test plan that gives positive measure to curves whose metric speed and the norm of the derivative are not equal.
@@
    There exists a function $f\in C^{1}(\mathbb{R}^{d})$ for which the minimal weak upper gradient and the metric slope (with respect to $\mathsf{d}_{g}$) are strictly larger than the norm of the gradient on a set of positive measure, i.e. $|Df|,|Df|_{w}>|\nabla_{g}f|_{g}$ on a set of positive $\mathrm{vol}_{g}$-measure.
4.2 A first example of non-infinitesimal Hilbertianity
In this subsection, we will more closely analyse the infinitesimal structure of a manifold similar to the previous example. This following example metric resembles the previous one, with the sole difference, that we only perturb the metric inside the interior of the unit cube $(0,1)^{d}\subset\mathbb{R}^{d}$ and leave it constant outside that set. We will prove that the resulting metric measure space is not infinitesimally Hilbertian.
@@ 4.2 A first example of non-infinitesimal Hilbertianity
$g:=2\mathbbm{1}_{((0,1)^{d})^{c}}Id+\mathbbm{1}_{(0,1)^{d}}\theta(x_{1})Id\in L^{\infty}(\mathbb{R}^{d},\mathbb{R}^{d\times d}).$
Note that $(\mathbb{R}^{d},g)$ satisfies (ii) from Assumption 3.1, $(\mathbb{R}^{d},\mathsf{d}_{g},\mathrm{vol}_{g})$ satisfies (2.1), and $g$ is constant outside the precompact set $(0,1)^{d}$. Inside $(0,1)^{d}$, $g$ the metric space arising from $(M,g)$ will behave just as in the previous subsection, we only restrict this behaviour to a compact set to conclude the lack of infinitesimal Hilbertianity.
@@
Proof of the claim. Indeed, choose a Lebesgue point $t$ of both $f^{-1}((-\infty,a])$ and $h^{\prime}$ and suppose for a contradiction that $h^{\prime}(t)\neq 0$. Then there exists a neighbourhood $U\ni t$ such that $h(s)\neq h(t)$ for every $s\in U\setminus\{t\}$. But as $t$ is a Lebesgue point of $f^{-1}((-\infty,a])$, we have that $(f^{-1}((-\infty,a])\cap U)\setminus\{t\}\neq\emptyset$. Hence, there exists a point in $s\in f^{-1}((-\infty,a])$ such that $h(s)\neq h(t)$, which contradicts the definition of $h$. As those Lebesgue points have full measure, this proves the claim.
@@
By Proposition 3.3, we get that for any curve $\gamma$ and $t\in[0,1]$ such that $\gamma_{t}\in O\times(0,1)^{d-1}\cup((0,1)^{d})^{c}$ and such that $\gamma$ is differentiable at $t$, it holds $|\dot{\gamma}_{t}|=|\dot{\gamma}_{t}|_{g(\gamma_{t})}$.
Next suppose that $y\in(0,1)^{d}$ is a Lebesgue point of $\{g=2\}$. In this argument, we will work with the definition of the metric $\mathsf{d}_{g}$ via (3.2). We will at first determine the shape and length of an almost distance minimising curve between two points that are close to each other. Fix an $\varepsilon>0$. Then, by the Lebesgue differentiation theorem, there exists a $\delta>0$ such that for each open interval $I\subset\mathbb{R}$ with $x_{1}(y)\in I$, and $\mathcal{L}^{1}(I)\leq\delta$, it holds
@@
Fix the null set $N^{(1)}_{g}\subset\mathbb{R}$ as the set of non-Lebesgue points of $\theta$ and define $N_{g}:=N^{(1)}_{g}\times\mathbb{R}^{d-1}$. We may assume that $O\cap N^{(1)}_{g}=\emptyset$.
Let $z\in(0,1)^{d}$ be such that $\mathsf{d}_{euc}(y,z)\leq\frac{1}{8}\mathrm{dist}_{euc}(y,\partial(0,1)^{d})$. Then any $2$-shortest curve connecting them, i.e. any curve $\gamma:[0,1]\to\mathbb{R}^{d}$ from $y$ to $z$ such that $\gamma\perp N_{g}$ and $L_{g}(\gamma)\leq 2\mathsf{d}_{g}(y,z)$, has to lie inside $(0,1)^{d}$. Let $\gamma:[0,1]\to(0,1)^{d}$ be a Lipschitz continuous curve transversal to $N_{g}$ such that $\gamma(0)=y$ and $\gamma(1)=z$. We will now estimate $\mathsf{d}_{g}(y,z)$ depending on $z-y\in\mathbb{R}^{d}$. We assume that $x_{1}(z)\geq x_{1}(y)$ and $|x_{1}(z)-x_{1}(y)|<\delta$. The case $x_{1}(z)\leq x_{1}(y)$ is analogous.
@@
$a_{z}-a_{y}\leq\delta,\ x_{1}(y)-a_{y}\leq\varepsilon,\ a_{z}-x_{1}(z)\leq\varepsilon,a_{z}-a_{y}\leq 2|z-y|_{euc}.$ (4.9)
@@
Then $\tilde{\gamma}$ is Lipschitz continuous and transversal to $N_{g}$. Indeed, on $(x_{1}\circ\gamma)^{-1}((-\infty,a_{y}]\cup[a_{z},\infty))$, we get that $\tilde{\gamma}\in\{a_{y},a_{z}\}\times(0,1)^{d-1}\subset N^{c}_{g}\cap\{g\equiv Id\}$ and on $(x_{1}\circ\gamma)^{-1}((a_{y},a_{z}))$, we have that $\gamma=\tilde{\gamma}$. Note that the methods from the proof of Lemma 4.2 yield that
@@
$=\int_{[0,1]}\sqrt{\theta(x_{1}(\tilde{\gamma}_{t}))}|\dot{\gamma}_{t}|_{euc}\,{\rm d}t=L_{g}(\tilde{\gamma}).$ (4.12)
@@
Step 2. Reduction to curves consisting of two straight lines.
@@
$L_{g}({\gamma})$ $=\int_{[0,1]}\sqrt{\theta(x_{1}(\tilde{\gamma}))}|\dot{\gamma}_{t}|_{euc}\,{\rm d}t$
@@
Write $O\cap(a_{y},a_{z}):=\bigcup_{n\in\mathbb{N}}I_{n}$, where $I_{n}$ are pairwise disjoint open intervals. By our choice of $\delta$ (see (4.8) and (4.9)), we get that $\mathcal{L}^{1}(O\cap(a_{y},a_{z}))\leq\varepsilon(a_{z}-a_{y})$. Now, by Lemma 4.2, we get the following estimate on the first component $(v_{(1)})_{1}$ of $v_{(1)}=\sum_{i=1}^{d}(v_{(1)})_{i}e_{i}$.
$|(v_{(1)})_{1}|=\Big{|}\sum_{n}\int_{(x_{1}\circ\gamma)^{-1}(I_{n})}\dot{(x_{1}\circ\gamma)}(t){\rm d}t\big{|}\leq\sum_{n}\mathcal{L}^{1}(I_{n})\leq\varepsilon|a_{z}-a_{y}|\leq 2\varepsilon|z-y|_{euc}.$
@@
$w_{(1)}:=v_{(1)}-(v_{(1)})_{1}e_{1}\in\{0\}\times\mathbb{R}^{d-1},\ w_{(2)}=v_{(2)}+(v_{(1)})_{1}e_{1}\in\mathbb{R}^{d}.$
@@
$L_{g}(\tilde{\gamma})\geq\sqrt{2}|w_{(2)}|_{euc}+|w_{(1)}|_{euc}-6\varepsilon|z-y|_{euc},$ (4.14)
where $w_{(1)}+w_{(2)}=z-y$ and $w_{(1)}\in\mathrm{span}(e_{2},\ldots e_{d})$. By potentially applying a linear transform from $\mathrm{SO}(d-1)$, we may assume that $z-y\in\mathrm{span}(e_{1},e_{2})$. We reduce the right hand side of (4.14), if we orthogonally project both $w_{(1)}$ and $w_{(2)}$ onto $\mathrm{span}(e_{1},e_{2})$, so we may assume that $w_{(1)},w_{(2)}\in\mathrm{span}(e_{1},e_{2})$. Now, if $z-y=ae_{1}$, for some $a\in\mathbb{R}$ we get that $w_{(1)}=0$ and hence
$L_{g}(\tilde{\gamma})\geq\sqrt{2}a-6a\varepsilon.$ (4.15)
Otherwise, we may write $z-y=a(be_{1}+e_{2})$, where $a,b\in\mathbb{R}$, $a\neq 0$. Write $w_{(1)}=a(1-c)e_{2}$ and $w_{(2)}=a(be_{1}+ce_{2})$. Now, we want to choose the optimal parameter $c\in[0,1]$ such that
@@
that is (4.14) without the error term, becomes minimal under the constraint that $w_{(1)}\in\mathrm{span}(e_{2})$. This reduces to finding $c$ such that
@@
More generally, writing $z-y=u=(u_{1},u^{\prime})$ where $u_{1}\in x_{1}(y)+(-\delta,\delta)\subset\mathbb{R}$ and $u^{\prime}\in\mathbb{R}^{d-1}$, we have that
@@
$\mathsf{d}_{g}(y,y+u)\geq\left\{\begin{array}[]{ll}|u_{1}|+|u^{\prime}|_{euc}-6\sigma(p,u_{1})|u|_{euc}&\mathrm{if}\ |u_{1}|\leq|u^{\prime}|_{euc},\\ \sqrt{2}|u|_{euc}-6\sigma(p,u_{1})|u|_{euc}&\mathrm{otherwise},\end{array}\right.$ (4.19)
@@
We will next prove an upper bound on $\mathsf{d}_{g}(y,z)$ and work with the regularisation. Fix $\epsilon>0$ For any $\varsigma>0$ and $g_{\varsigma}:=\rho_{\varsigma}*g$, it holds $g_{\varsigma}\leq 2Id$. Hence, for all $\varsigma>0$, we have that
@@
Moreover, we can find a rational $q\in\mathbb{Q}\cap(0,1)$ such that $|q-x_{1}(y)|\leq\epsilon|u|_{euc}$. Fix a $10<\alpha<\beta<$ 1 such that $U:=(\alpha,\beta)^{d-1}$ satisfies that $y,z\in(0,1)\times U$. Now for some $\varsigma_{\epsilon}>0$ small enough, we have that for $\varsigma\in(0,\varsigma_{\epsilon})$, it holds $g_{\varsigma}=Id$ on $\{q\}\times U$. If $|u_{1}|<|u^{\prime}|_{euc}$, write $b=\frac{u_{1}}{|u^{\prime}|_{euc}}$. Then define the curves
@@
Moreover as $g_{\varsigma}\equiv Id$ on $\{q\}\times U\supset\gamma_{2}([0,1])$, we get that
@@
$0\displaystyle 0\leq\lim_{h\to 0}\frac{\alpha(hu)-\mathsf{d}_{g}(\gamma^{y,u}_{0},\gamma^{y,u}_{h})}{|h|}\leq\lim_{h\to 0}\frac{\sigma(y,hu_{1})|hu|_{euc}}{|h|}=\lim_{h\to 0}\sigma(y,hu_{1})|u|_{euc}=$ 0 .

### v3

@@ 3 The distance on manifolds with non-continuous Riemannian metrics
We may assume to be working in $\mathbb{R}^{d}$. Now there exist $\lambda,\Lambda>0$ such that $\lambda\mathrm{Id}\leq g\leq\Lambda\mathrm{Id}$ on $K$. Denote $v:=\dot{\gamma}^{1}_{t}$ and $w:=\dot{\gamma}^{2}_{t}$. By translation we may assume that $t=0$ and denote $p:=\gamma^{1}_{0}=\gamma^{2}_{0}$. By the previous proposition, we may assume that $\gamma^{1}_{t}=p+tv$ and $\gamma^{2}_{t}=p+tw$. But then
@@
Given a manifold $M$, a distance $\mathsf{d}$ on $M$, and a measurable positive definite inner product $g$ on $TM$, we call $(M,\mathsf{d},g)$ a quasi-Riemannian manifold if there exists an atlas $\mathcal{A}=\{W,\psi\}$ such that the following conditions hold:
@@
    In each $(W,\psi)$, there exist constants $\lambda,\Lambda>0$ such that for any $x,y\in W$, it holds
    $\lambda\leq(\psi_{*}g)^{ij}\partial_{x^{i}}\mathsf{d}^{W}_{euc}(x,y)\partial_{x^{j}}\mathsf{d}^{W}_{euc}(x,y)\leq\Lambda.$
@@
    $\limsup_{y\to x}\frac{|f(x)-f(y)|}{\mathsf{d}(x,y)}=|\nabla_{g}f|_{g}(x).$
@@ 4 Construction of examples
In this section, we will construct Riemannian manifolds for with non-continuous metrics that together with the distance $\mathsf{d}_{g}$ as defined in the previous section, do not give rise to quasi-Riemannian manifolds or infinitesimally Hilbertian metric measure spaces.
4.1 An introductory example of a non-quasi Riemannian manifold
We will now construct a metric $g$ on $\mathbb{R}^{d}$ for $d\geq 2$ such that $g,g^{-1}\in L^{\infty}$ and such that there exists a smooth function $f:\mathbb{R}^{d}\to\mathbb{R}$ such that $|Df|_{w},|Df|>|\nabla_{g}f|_{g}$ on a set of positive measure, which implies that $(M,\mathsf{d}_{g},g)$, where $\mathsf{d}_{g}$ is defined as in (3.2), is not a quasi-Riemannian manifold. In [37], Sturm proved that given a Riemannian manifold $(M,g)$, $g,g^{-1}\in L^{\infty}_{\rm loc}$, there exists a metric $g^{\prime}$ on $M$ such that $^{\prime},{g^{\prime}}^{-1}\in L^{\infty}_{\rm loc}, g<g$$^{\prime}<g$ on a set of positive $\mathrm{vol}_{g}$-measure and $\mathsf{d}_{g}=\mathsf{d}_{g^{\prime}}$, which immediately shows that Questions 1 and 2 may have negative answers. The example presented in this subsection is related to the one in [10] and specifically designed to pave the way for examples, where also infinitesimal Hilbertianity fails, that will be discussed in later subsections
@@ 4.1 An introductory example of a non-quasi Riemannian manifold
$g(x_{1},\ldots,x_{d}):=\theta(x_{1})\mathrm{Id}\in L^{\infty}(\mathbb{R}^{d},\mathbb{R}^{d\times d}).$
Then $(\mathbb{R}^{d},g)$ satisfies Assumption 3.1 and $(\mathbb{R}^{d},\mathsf{d}_{g},\mathrm{vol}_{g})$ satisfies condition (2.1). We observe that
@@
Proof of the claim. Fix $t,h$ as above and fix an $\varepsilon>0$. Note that from $g\geq\mathrm{Id}$, we get that $\mathsf{d}_{g}\geq\mathsf{d}_{euc}$, hence $\mathsf{d}_{g}(\gamma^{y}_{t},\gamma^{y}_{t+h})\geq h$. Thus, it remains to prove that $\mathsf{d}_{g}(\gamma^{y}_{t},\gamma^{y}_{t+h})\leq h$. There exists a $q\in\mathbb{Q}\cap(0,1)$ such that $|x_{1}(y)-q|\leq\varepsilon$. Let $i\in\mathbb{N}$ be such that $q_{i}=q$ in the chosen counting. Then define the curves
@@
Denote by $\lambda$ the concatenation of those three curves and note that it yields a piecewise smooth curve that connects $\gamma^{y}_{t}$ and $\gamma^{y}_{t+h}$. Now, for all $\varsigma>0$, it holds $g_{\varsigma}=g*\rho_{\varsigma}\leq 2\mathrm{Id}$, so we get that $L_{g_{\varsigma}}(\lambda_{1})+L_{g_{\varsigma}}(\lambda_{3})\leq 4\varepsilon$. For $\varsigma\leq\frac{\kappa}{2^{i+2}}$ that $g_{\varsigma}|_{\{x_{1}=q\}}\equiv\mathrm{Id}$, hence $L_{g_{\varsigma}}(\lambda_{2})=h$. Thus,
@@
Note that this is indeed a feasible test plan, as ${\rm d}\mathcal{L}^{d}\leq{\rm d}\mathrm{vol}_{g}\leq 2^{\frac{d}{2}}{\rm d}\mathcal{L}^{d}$. But now for all $y\in E$, almost all $t\in[0,1]$, we get that $g(\gamma^{y}_{t})\geq\frac{3}{2}\mathrm{Id}$ and $\dot{\gamma}^{y}_{t}=e_{2}$ and $|\dot{\gamma}^{y}_{t}|_{g}\geq\sqrt{\frac{3}{2}}$. But by (4.3), we get that the metric speed $|\dot{\gamma}^{y}_{t}|=1$ for all $y,t$, hence we found a test plan that gives positive measure to curves whose metric speed and the norm of the derivative are not equal.
@@
    There exists a function $f\in C^{1}(\mathbb{R}^{d})$ for which the minimal weak upper gradient and the metric slope are strictly larger than the norm of the gradient on a set of positive measure, i.e., $|Df|,|Df|_{w}>|\nabla_{g}f|_{g}$ on a set of positive $\mathrm{vol}_{g}$-measure.
4.2 An introductory example of non-infinitesimal Hilbertianity
In this subsection, we will more closely analyse the infinitesimal structure of a manifold similar to the previous example. This following example metric resembles the previous one, with the sole difference, that we only perturb the metric inside the interior of the unit cube $(0,1)^{d}\subset\mathbb{R}^{d}$ for $d\geq 2$ and leave it constant outside that set. We will prove that the resulting metric measure space is not infinitesimally Hilbertian.
@@ 4.2 An introductory example of non-infinitesimal Hilbertianity
$g(x_{1},\ldots,x_{d}):=2\mathbbm{1}_{((0,1)^{d})^{c}}\mathrm{Id}+\mathbbm{1}_{(0,1)^{d}}\theta(x_{1})\mathrm{Id}\in L^{\infty}(\mathbb{R}^{d},\mathbb{R}^{d\times d}).$
Note that $(\mathbb{R}^{d},g)$ satisfies Assumption 3.1, $(\mathbb{R}^{d},\mathsf{d}_{g},\mathrm{vol}_{g})$ satisfies (2.1), and $g$ is constant outside the precompact set $(0,1)^{d}$. Inside $(0,1)^{d}$, $g$ the metric space arising from $(M,g)$ will behave just as in the previous subsection, we only restrict this behaviour to a compact set to conclude the lack of infinitesimal Hilbertianity.
@@
Proof of the claim. Indeed, choose a point $t\in[0,1]$ that is a differentiability point of $h$ and that is a Lebesgue point of both $f^{-1}((-\infty,a])$ and $h^{\prime}$. Suppose for a contradiction that $h^{\prime}(t)\neq 0$. Then there exists a neighbourhood $U\ni t$ such that $h(s)\neq h(t)$ for every $s\in U\setminus\{t\}$. But as $t$ is a Lebesgue point of $f^{-1}((-\infty,a])$, we have that $(f^{-1}((-\infty,a])\cap U)\setminus\{t\}\neq\emptyset$. Hence, there exists a point in $s\in f^{-1}((-\infty,a])$ such that $h(s)\neq h(t)$, which contradicts the definition of $h$. As those Lebesgue points have full measure, this proves the claim.
@@
With Proposition 3.3, we will now determine the metric speed of a curve through a point $y\in\mathbb{R}^{d}$ depending on the derivative.
First suppose that $y\in(0,1)^{d}$ and that $y$ is a Lebesgue point of $\{g=2\}$. In this argument, we will work with the definition of the metric $\mathsf{d}_{g}$ via (3.2). We will at first determine the shape and length of an almost distance minimising curve between two points that are close to each other. Fix an $\varepsilon>0$. Then, by the Lebesgue differentiation theorem, there exists a $\delta>0$ such that for each open interval $I\subset\mathbb{R}$ with $x_{1}(y)\in I$, and $\mathcal{L}^{1}(I)\leq\delta$, it holds
@@
Fix the null set $N^{(1)}_{g}\subset\mathbb{R}$ as the set of non-Lebesgue points of $\theta$ and define $N_{g}:=N^{(1)}_{g}\times\mathbb{R}^{d-1}$. Note that then $O\cap N^{(1)}_{g}$$\emptyset$.
Let $z\in(0,1)^{d}$ be such that $\mathsf{d}_{euc}(y,z)\leq\frac{1}{8}\mathrm{dist}_{euc}(y,\partial(0,1)^{d})$. Then any $2$-shortest curve connecting them, i.e., any curve $\gamma:[0,1]\to\mathbb{R}^{d}$ from $y$ to $z$ such that $\gamma\perp N_{g}$ and $L_{g}(\gamma)\leq 2\mathsf{d}_{g}(y,z)$, has to lie inside $(0,1)^{d}$. Let $\gamma:[0,1]\to(0,1)^{d}$ be a Lipschitz continuous curve transversal to $N_{g}$ such that $\gamma(0)=y$ and $\gamma(1)=z$. We will now estimate $\mathsf{d}_{g}(y,z)$ depending on $z-y\in\mathbb{R}^{d}$. We assume that $x_{1}(z)\geq x_{1}(y)$ and $|x_{1}(z)-x_{1}(y)|<\delta$. The case $x_{1}(z)\leq x_{1}(y)$ is analogous.
@@
$a_{z}-a_{y}\leq\delta,\quad x_{1}(y)-a_{y}\leq\varepsilon,\quad a_{z}-x_{1}(z)\leq\varepsilon,\quad a_{z}-a_{y}\leq 2|z-y|_{euc}.$ (4.9)
@@
Then $\tilde{\gamma}$ is Lipschitz continuous and transversal to $N_{g}$. Indeed, on $(x_{1}\circ\gamma)^{-1}((-\infty,a_{y}]\cup[a_{z},\infty))$, we get that $\tilde{\gamma}\in\{a_{y},a_{z}\}\times(0,1)^{d-1}\subset N^{c}_{g}\cap\{g\equiv\mathrm{Id}\}$ and on $(x_{1}\circ\gamma)^{-1}((a_{y},a_{z}))$, we have that $\gamma=\tilde{\gamma}$. Note that the methods from the proof of Lemma 4.2 yield that
@@
[…]
@@
Step 2. Reduction to the length of curves consisting of two straight lines.
@@
Write $O\cap(a_{y},a_{z}):=\bigcup_{n\in\mathbb{N}}I_{n}$, where $I_{n}$ are pairwise disjoint open intervals. By our choice of $\delta$ (see (4.8) and (4.9)), we get that $\mathcal{L}^{1}(O\cap(a_{y},a_{z}))\leq\varepsilon(a_{z}-a_{y})$. Now, by Lemma 4.2, we get the following estimate on the first component $x_{1}(v_{(1)})$ of $v_{(1)}=\sum_{i=1}^{d}x_{i}(v_{(1)})e_{i}$.
$|x_{1}(v_{(1)})|=\Big|\sum_{n}\int_{(x_{1}\circ\gamma)^{-1}(I_{n})}\dot{(x_{1}\circ\gamma)}(t){\rm d}t\big|\leq\sum_{n}\mathcal{L}^{1}(I_{n})\leq\varepsilon|a_{z}-a_{y}|\leq 2\varepsilon|z-y|_{euc}.$
@@
$w_{(1)}:=v_{(1)}-x_{1}(v_{(1)})e_{1}\in\{0\}\times\mathbb{R}^{d-1},\ w_{(2)}=v_{(2)}+x_{1}(v_{(1)})e_{1}\in\mathbb{R}^{d}.$
@@
$L_{g}({\gamma})\geq\sqrt{2}|w_{(2)}|_{euc}+|w_{(1)}|_{euc}-6\varepsilon|z-y|_{euc},$ (4.14)
where $w_{(1)}+w_{(2)}=z-y$ and $w_{(1)}\in\mathrm{span}(e_{2},\ldots e_{d})$.
Step 3. Optimisation of $w_{(1)}$ and $w_{(2)}$ under the above constraints.
By potentially applying a linear transform from $\mathrm{SO}(d-1)$ that rotates the last d1d-1 entries, we may assume that $z-y\in\mathrm{span}(e_{1},e_{2})$. We reduce the right hand side of (4.14), if we orthogonally project both $w_{(1)}$ and $w_{(2)}$ onto $\mathrm{span}(e_{1},e_{2})$, so we may assume that $w_{(1)},w_{(2)}\in\mathrm{span}(e_{1},e_{2})$. Now, if $z-y=ae_{1}$, for some $a\in\mathbb{R}$ we get that $w_{(1)}=0$ and hence
$L_{g}({\gamma})\geq\sqrt{2}a-6a\varepsilon.$ (4.15)
Otherwise, we may write $z-y=a(be_{1}+e_{2})$, where $a,b\in\mathbb{R}$, $a\neq 0$. Write $w_{(1)}=a(1-c)e_{2}$ and $w_{(2)}=a(be_{1}+ce_{2})$. Now, we want to choose the optimal parameter $c\in\mathbb{R}$ such that
@@
that is (4.14) without the error term, becomes minimal under the constraint that $w_{(1)}\in\mathrm{span}(e_{2})$. The error term in (4.14) is independent of $c$, hence this reduces to finding $c\in\mathbb{R}$ such that
@@
Step 4. Description of the metric speed of a curve depending on its derivative $u$.
Recall that we chose $y$ to be a Lebesgue point of $\{g=2\}$. Writing $z-y=u=(u_{1},u^{\prime})$ where $u_{1}\in x_{1}(y)+(-\delta,\delta)\subset\mathbb{R}$ and $u^{\prime}\in\mathbb{R}^{d-1}$, the previous three steps yield that
@@
$\mathsf{d}_{g}(y,y+u)\geq\left\{\begin{array}[]{ll}|u_{1}|+|u^{\prime}|_{euc}-6\sigma(y,u_{1})|u|_{euc}&\mathrm{if}\ |u_{1}|\leq|u^{\prime}|_{euc},\\ \sqrt{2}|u|_{euc}-6\sigma(y,u_{1})|u|_{euc}&\mathrm{otherwise},\end{array}\right$. (4.19)
@@
We will next prove an upper bound on $\mathsf{d}_{g}(y,z)$ by working with the regularisation, i.e., smooth metrics $g_{\epsilon}$ that converge to $g$ almost everywhere and such that $\mathsf{d}_{g_{\epsilon}}\to\mathsf{d}_{g}$. We will construct curves $\hat{\gamma}$ that are given by the concatenation of four straight lines, two of which will be arbitrarily short. The remaining two lines will be optimal in the sense of steps 2 and 3. Fix $\epsilon>0$. For any $\varsigma>0$ and $g_{\varsigma}:=\rho_{\varsigma}*g$, it holds $g_{\varsigma}\leq 2\mathrm{Id}$. Hence, for all $\varsigma>0$, we have that
@@
Moreover, we can find a rational $q\in\mathbb{Q}\cap(0,1)$ such that $|q-x_{1}(y)|\leq\epsilon|u|_{euc}$. Fix a $0<\alpha<\beta<1$ such that $U:=(\alpha,\beta)^{d-1}$ satisfies that $y,z\in(0,1)\times U$. Now for some $\varsigma_{\epsilon}>0$ small enough, we have that for $\varsigma\in(0,\varsigma_{\epsilon})$, it holds $g_{\varsigma}=\mathrm{Id}$ on $\{q\}\times U$. If $|u_{1}|<|u^{\prime}|_{euc}$, write $b=\frac{u_{1}}{|u^{\prime}|_{euc}}$. Then define the curves
@@
Moreover as $g_{\varsigma}\equiv\mathrm{Id}$ on $\{q\}\times U\supset\gamma_{2}([0,1])$, we get that
@@
$0\leq\lim_{h\to 0}\frac{\alpha(hu)-\mathsf{d}_{g}(\gamma^{y,u}_{0},\gamma^{y,u}_{h})}{|h|}\leq\lim_{h\to 0}\frac{\sigma(y,hu_{1})|hu|_{euc}}{|h|}=\lim_{h\to 0}\sigma(y,hu_{1})|u|_{euc}=0.$